Flamo:

Contestants in a long distance race run 5 kilometers due south. The runners then turn right and head west towards the finish line. If the straight line distance from start to finish is 13 kilometers, what is the distance that the contestants traveled while heading west?

Flamo:

So Same thing?

Angle:

mhm!

Flamo:

Okay.

Flamo:

a^2 + b^2 = c^2?

Angle:



Flamo:

o

Flamo:

13^2 + 5^2 = c^2?

Angle:

not quite they are saying the hypotenuse is 13

Flamo:

o, Woops XD

Flamo:

a^2 + 5^2 = 15^2

Angle:

yup!

Flamo:

No, Not 15 13... -.-

Angle:

lol

Flamo:

a^2 + 5^2 = 13^2

Flamo:

a^2 + 25 = 169

Flamo:

a^2 = 169 - 25?

Angle:

yes, that's good :)

Flamo:

Ye!

Flamo:

144 to the Square root is 12

Flamo:

So 12 Km :D

Angle:

great job!

Flamo:

Ye!

Flamo:

1 More Question That I need help on.

Angle:

okie